The length of a rectangle is 3 inches more than the width. The perimeter of the rectangle is 24 inches......How do I solve this??? HELP!!
Answer by checkley71(8403)

L=W+3
24=2L+2W
24=2(W+3)+2W
24=2W+6+2W
24=4W+6
4W=24-6
4W=18
W=18/4
W=4.5 ANSWER.
L=4.5+3
L=7.5 ANSWER.
PROOF
24=2*4.5+2*7.5
24=9+15
24=24